# ICPO: Intrinsic Confidence-Driven Group Relative Preference Optimization for Efficient Reinforcement Learning

### v4

@@ Appendix C Theoretical Explanation
We systematically explain the effectiveness of ICPO from three aspects: (1) why ICPO can calibrate policy updates; (2) why ICPO can address the coarse-grained reward problem; and (3) why ICPO can mitigate the interference from noisy rewards.
C.1 Aspect 1: Calibrate Policy Updates
@@ 5. Calculate the relative advantages of different responses in vanilla GRPO and ICPO:
C.2 Aspect 2: Tackle Coarse-grained Reward
The coarse-grained reward problem refers to the scenario where, under the same prompt, all generated responses receive identical rewards (typically stemming from the coarse-grained design of the reward function). In such cases, GRPO is unable to compute the relative advantages among responses within the group, resulting in a lack of effective gradient signals for policy optimization. This, in turn, triggers policy collapse and a sharp decline in model performance.
@@ C.2 Aspect 2: Tackle Coarse-grained Reward
C.3 Aspect 3: Mitigate Noisy Reward

### v5

@@ Appendix E Theoretical Explanation
We systematically explain the effectiveness of ICPO from three aspects: (1) why ICPO can calibrate policy updates; (2) why ICPO can address the sparse reward problem; and (3) why ICPO can mitigate the interference from noisy rewards.
E.1 Aspect 1: Calibrate Policy Updates
@@ 5. Calculate the relative advantages of different responses in vanilla GRPO and ICPO:
E.2 Aspect 2: Tackle sparse Reward
The sparse reward problem refers to the scenario where, under the same prompt, all generated responses receive identical rewards (typically stemming from the sparse design of the reward function). In such cases, GRPO is unable to compute the relative advantages among responses within the group, resulting in a lack of effective gradient signals for policy optimization. This, in turn, triggers policy collapse and a sharp decline in model performance.
@@ E.2 Aspect 2: Tackle sparse Reward
E.3 Aspect 3: Mitigate Noisy Reward